OpenStudy (anonymous):

What is the derivative of y=e^(e^(sin2x))??

OpenStudy (anonymous):

y' = e^(e^(sin 2x)) * e^(sin 2x) * cos 2x * 2 y' = 2*(cos 2x)*e^(e^(sin 2x) + sin 2x)